OpenStudy (anonymous):

As part of a fundraiser, a local charity has sold 10 raffel tickets, with a goal of selling 60 tickets. What percentage of the goal has been sold? round to the nearest tenth of a percent

OpenStudy (anonymous):

10/60 *100

OpenStudy (nowhereman):

yes, but don't forget the % sign. It is just another way to say 1/100. \[\text{percentage} = \frac{\text{amount}}{\text{total}}\cdot 100\%\]

OpenStudy (anonymous):

whats the answer?

OpenStudy (nowhereman):

Don't you have a calculator?

OpenStudy (anonymous):

im saying do you do 10 divided by 60

OpenStudy (nowhereman):

yeah, thats what ankur said, didn't he?

OpenStudy (anonymous):

yeah i was just making sure that was correct